OpenStudy (anonymous):

in integrating by parts (calculus) .. can anyone give me steps on how to , because most of the time i don't know what to do next.

OpenStudy (anonymous):

\[\int\limits u dv=uv- \int\limits v du\] you will need to choose u (something to differentiate) and dv (something to integrate)(easy) example: \[\int\limits\limits x sinx dx\] u=x dv=sinx du=dx v=-cosx \[\int\limits\limits x sinx dx\]=\[-x cosx- \int\limits -cosx dx=-xcosx+\int\limits cosx\] =-x cosx +sinx +c